OpenStudy (anonymous):

The sum of four consecutive integers is 74. What is the first integer? 16 17 18 19

OpenStudy (texaschic101):

x + x + 1 + x + 2 + x + 3 = 74 now combine like terms

OpenStudy (anonymous):

let four consecutive number be n,n+1,n+2,n+3 so n+n+1+n+2+n+3=74 4n+6=74 4n=68 =>n=17

OpenStudy (texaschic101):

well...there is you answer

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (solidicecream):

. .